# Strong phonon-mediated high temperature superconductivity in Li2AuH6 under ambient pressure

### v1

@@ Abstract
Introduction Searching superconducting materials with high transition temperature () is a hot issue since the discovery of superconductivity [1]. As the lightest element, high- superconductivity was anticipated to realize in metallic hydrogen [2]. However, solid hydrogen is too difficult to stabilize and show superconductivity under ambient pressure. Consequently, researchers are shifting their focus toward hydrogen-rich compounds under high-pressure conditions. In 2014, H3S was theoretically predicted to exhibit superconductivity under high pressure [3, 4], which was experimentally confirmed [5]. With development of high-pressure technique, various hydrogen-rich superconductors with high- were synthesized, for example, LaH10 [6, 7], YH9 [8, 9], and CaH6 [10], which makes high-pressure hydride become a very promising candidate for room-temperature superconductivity.
@@
In recent years, ternary hydride gets more attentions due to the enlarged phase space and enhanced flexibility of manipulation. Some superconducting ternary hydrides were theoretically proposed, for example, H6SX (X = Cl, Br) [18], LaRH8 (R = B, Be) [19, 20, 21], Li2MgH16 [22], KB2H8 [23], and CsBH5 [24]. Among them, the strong electron-phonon coupling (EPC) driven by metallic B-H -bond and related phonon in KB2H8 lead a superconducting transition with 134 K under 12 GPa. Furthermore, using chemical doping to modulate virtual high-pressure effect, CsBH5 exhibit superconductivity with 98 K under near-ambient 1 GPa pressure. These works suggest that ternary hydride is expected to exhibit high- superconductivity under relatively low even ambient pressure.
@@
In this letter, we performed vast investigations regarding ternary hydrides through our developed AI search engine, namely InvDesFlow [28, 29], which integrates generative AI and several graph neural networks for the discovery of high- superconductors. A high- superconductor Li2AuH6 that is isostructural to Mg2IrH6 was identified. Combined with the first-principles density-functional theory and superconducting EPC calculations, Li2AuH6 shows a high- 140 K under ambient pressure. Thermodynamic study suggests that Li2AuH6 is a metastable phase and may be experimentally synthesized via the known LiH and LiAu compounds. Detailed EPC analysis indicates that the H-1 electrons couple with special phonon modes then drive strong Cooper pairing. Different from previous reports that emphasized the vibrations of H octahedrons, we find that the vibrations of Li atoms also contribute strong EPC, even stronger than that of vibrations of H octahedrons. Our work reveals a ternary hydride Li2AuH6 with superconducting 140 K under ambient pressure and suggests that EPC can be enhanced by introducing more strong coupling modes.
Methods In this work, the DFT calculations were performed by the first-principles package QUANTUM-ESPRESSO [30]. The generalized gradient approximation of the Perdew-Burke-Ernzerhof formula [31] for the exchange and correlation functional was chosen. The optimized norm-conserving Vanderbilt pseudopotential was used [32]. The kinetic energy cutoff and the charge density cutoff were respectively set to be 80 Ry and 320 Ry. The charge densities were self-consistently calculated on an unshifted 161616 -point mesh grid with a Methfessel-Paxton smearing of 0.02 Ry [33]. And the phonon were calculated on a 444 -point mesh grid based on the density-functional perturbation theory [34].
@@
Hence, we propose a feasible route 6LiH + 5Au Li2AuH6 + 4LiAu. The calculated formation energy of right side of equation is only 0.038 eV/atom higher than that of left side. According to a survey [41], the lower formation energy than a typical value of 80 meV/atom implies a good experimental synthesizability. These results suggest that Li2AuH6 is a metastable phase that may be experimentally synthesized.
@@
Discussion and Conclusion The establishment of Bardeen-Cooper-Schrieffer (BCS) theory [43] provides a theoretical support for understanding and searching conventional superconductors. And metallizing the -bonding electrons [44, 45] further proposes a possible way in finding high- superconductors, which has also achieved success in explaining superconductivity of MgB2 [46]. Specifically, the -bonds formed by the in-plane B-2 electrons strongly couple with an in-plane vibration mode of boron atoms [47]. Guiding by this, lots of theoretical works are focusing on finding materials with metallic -bands. B possesses close electronegativity to H, hence B-H system is regarded as a promising candidate. For example, KB2H8 is predicted to be a high- BCS superconductor due to the strong EPC between the metallic B-H -electrons and H-related phonon modes [23]. However, our studies here find the Au-H system without metallic -electrons can also exhibit high- superconductivity, which leads us to realize that the strong coupling phonon that can cause a significant change of electronic density may be a more critical factor.
Hence, we propose a concept of BCS superconducting unit, which provides strong EPC phonon modes as superconducting pairing glue, just like the superconducting unit of CuO planes provides antiferromagnetic fluctuation as glue for Cooper pairing in cuprates. Furthermore, H octahedron with strong EPC phonon modes may be a potential BCS superconducting unit. One can try to design or synthesize more similar BCS superconducting units, such as H tetrahedron, etc., and stabilize them in crystal lattice to find high- BCS superconductors. In addition, considering the rising approaches of high-throughput calculations, crystal structure prediction, and artificial intelligence, intercalating atoms directly in existing materials may introduce strongly coupled phonons into the system, which is a feasible way to investigate superconducting multicomponent compounds.
@@
This work was financially supported by the National Natural Science Foundation of China (Grant No.62476278, No.12434009, and No.12204533). Z.Y.L. was also supported by Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302402). Computational resources have been provided by the Physical Laboratory of High Performance Computing at Renmin University of China.

### v2

# High temperature superconductivity in Li2AuH6 mediated by strong electron-phonon coupling under ambient pressure

@@ Abstract
pacs:
Introduction. Searching superconducting materials with high transition temperature () is a hot issue since the discovery of superconductivity [1]. As the lightest element, high- superconductivity was anticipated to realize in metallic hydrogen [2]. However, solid hydrogen is too difficult to stabilize and show superconductivity under ambient pressure. Consequently, researchers are shifting their focus toward hydrogen-rich compounds under high-pressure conditions. In 2014, H3S was theoretically predicted to exhibit superconductivity under high pressure [3, 4], which was experimentally confirmed [5]. With development of high-pressure technique, various hydrogen-rich superconductors with high- were synthesized, for example, LaH10 [6, 7], YH9 [8, 9], and CaH6 [10], which makes high-pressure hydride become a very promising candidate for room-temperature superconductivity.
@@
In recent years, ternary hydride gets more attentions due to the enlarged phase space and enhanced flexibility of manipulation. Some superconducting ternary hydrides were theoretically proposed, for example, H6SX (X = Cl, Br) [18], LaRH8 (R = B, Be) [19, 20, 21], Li2MgH16 [22], KB2H8 [23], and CsBH5 [24]. Among them, the strong electron-phonon coupling (EPC) driven by metallic B-H -bond and related phonon in KB2H8 lead a superconducting transition with 134 K under 12 GPa. Furthermore, using chemical doping to modulate virtual high-pressure effect, CsBH5 exhibit superconductivity with 98 K under near-ambient 1 GPa pressure. In addition, ternary hydrides LaBeH8 [25] ( 110 K under 80 GPa), LaB2H8 [26] ( 106 K under 90 GPa), and La-Ce-H system [27] ( 176 K under 100 GPa) were experimentally reported. Both of these theoretical and experimental works suggest that ternary hydride is expected to exhibit high- superconductivity under relatively low even ambient pressure.
@@
Techniques from machine learning and data science are increasingly being employed to address problems in materials science [31]. In particular, generative modeling methods based on autoencoder architectures and diffusion [32, 33] have been employed to generate crystal structures. In this letter, we performed vast investigations regarding ternary hydrides through our developed AI search engine, namely InvDesFlow [32], which integrates generative AI and several graph neural networks for the discovery of high- superconductors. See the supplementary materials [34] for the detail of the mentioned AI search engine (see also references [2, 9] therein). The specific setting of the AI model can be found in the source code [35]. A high- superconductor Li2AuH6 that is isostructural to Mg2IrH6 was identified. Combined with the first-principles density-functional theory and superconducting EPC calculations, Li2AuH6 shows a high- 140 K under ambient pressure. Thermodynamic study suggests that Li2AuH6 is a metastable phase and may be experimentally synthesized via the known LiH and LiAu compounds. Detailed EPC analysis indicates that the H-1 electrons couple with special phonon modes then drive strong Cooper pairing. Different from previous reports that emphasized the vibrations of H octahedrons, we find that the vibrations of Li atoms also contribute strong EPC, even stronger than that of vibrations of H octahedrons. Our work reveals a ternary hydride Li2AuH6 with superconducting 140 K under ambient pressure and suggests that EPC can be enhanced by introducing more strong coupling modes.
Methods. In this work, the DFT calculations were performed by the first-principles package QUANTUM-ESPRESSO [36]. The generalized gradient approximation of the Perdew-Burke-Ernzerhof formula [37] for the exchange and correlation functional was chosen. The optimized norm-conserving Vanderbilt pseudopotential was used [38]. The kinetic energy cutoff and the charge density cutoff were respectively set to be 80 Ry and 320 Ry. The charge densities were self-consistently calculated on an unshifted 161616 -point mesh grid with a Methfessel-Paxton smearing of 0.02 Ry [39]. And the phonon were calculated on a 444 -point mesh grid based on the density-functional perturbation theory [40].
@@
Hence, we propose a feasible route 6LiH + 5Au Li2AuH6 + 4LiAu. The calculated formation energy of right side of equation is only 38 meV/atom higher than that of left side. As for such a solid metastable phase, the reaction kinetics under ambient condition may be slow, hence catalysts may be used. In addition, metastable structure may decompose or transform into other phases during synthesis, which suggests that rapid quenching or inert gas atmospheres or encapsulation may be needed. In a word, according to a survey [47], the lower formation energy than a typical value of 80 meV/atom implies a good experimental synthesizability. These results suggest that Li2AuH6 is a metastable phase that may be experimentally synthesized.
@@
Discussion and Conclusion. The establishment of Bardeen-Cooper-Schrieffer (BCS) theory [49] provides a theoretical support for understanding and searching conventional superconductors. And metallizing the -bonding electrons [50, 51] further proposes a possible way in finding high- superconductors, which has also achieved success in explaining superconductivity of MgB2 [52]. Specifically, the -bonds formed by the in-plane B-2 electrons strongly couple with an in-plane vibration mode of boron atoms [53]. Guiding by this, lots of theoretical works are focusing on finding materials with metallic -bands. B possesses close electronegativity to H, hence B-H system is regarded as a promising candidate. For example, KB2H8 is predicted to be a high- BCS superconductor due to the strong EPC between the metallic B-H -electrons and H-related phonon modes [23]. However, our studies here find the Au-H system without metallic -electrons can also exhibit high- superconductivity, which leads us to realize that the strong coupling phonon that can cause a significant change of electronic density may be a more critical factor.
Hence, we propose a concept of BCS superconducting unit, which provides strong EPC phonon modes as superconducting pairing glue, just like the superconducting unit of CuO planes provides antiferromagnetic fluctuation as glue for Cooper pairing in cuprates. Furthermore, H octahedron with strong EPC phonon modes may be a potential BCS superconducting unit. One can try to design or synthesize more similar BCS superconducting units, such as H tetrahedron, etc., and stabilize them in crystal lattice to find high- BCS superconductors. In addition, considering the rising approaches of high-throughput calculations, crystal structure prediction, and artificial intelligence (AI), intercalating atoms directly in existing materials may introduce strongly coupled phonons into the system, which is a feasible way to investigate superconducting multicomponent compounds.
In this work, we first employed advanced AI technologies [32] to identify candidate materials with high , and then conducted high-precision DFT calculations for the candidate materials recommended by AI. This approach differs from the traditional high-throughput DFT screening of materials. High-throughput methods, limited by the computational cost of DFT, can only perform coarse calculations for all materials, which frequently leads to misleading results due to insufficient computational accuracy. In contrast, AI-based searching can more precisely narrow down the range of target materials, thereby allowing more computational resources (such as introducing anisotropic corrections, refined electron-phonon coupling analysis, and other advanced computational strategies) to be allocated to more valuable candidate materials.
@@
Acknowledgments. This work was financially supported by the National Natural Science Foundation of China (Grant No.62476278, No.12434009, and No.12204533). Z.Y.L. was also supported by the National Key R&D Program of China (Grants No. 2024YFA1408601). Computational resources have been provided by the Physical Laboratory of High Performance Computing at Renmin University of China.